If 10ml of 0.1M, CH3COOH and 10ml of 0.1M, CH3COONa are mixed the pH of the resultant buffer solution is _______(Ka = 1.82x10-5).
1
Expert's answer
2020-09-17T07:16:37-0400

From Henderson-Hasselbalch equation:

pH=pKa+logCconj.base/Cacid

pKa=-logKa=-log1.85×10-5=4.74.

While mixing of equal volumes of two solutions the resulting volume is two times higher than each one of the initial solutions, so the resulting concentrations are two times lower:

C(CH3COOH)=C(CH3COONa)=0.1/2=0.05(mol/L).

pH=4.74+log 0.05/0.05=4.74+0=4.74.